# Stochastic Online Metric Matching: Adversarial is no Harder than Stochastic

### v1

@@ Appendix B Cost Functions Beyond Triangle Inequality
The proof of Theorem 7 can be obtained by replacing every usage of the triangle inequality by the approximate triangle inequality in the proof of Theorem 1 and hence is omitted.
Next, we consider the case where $X=[0,1]^{d}$ and $\delta(x,y)=\left\|{x-y}\right\|_{2}^{p}$ for $p\geq 1$, which we will prove that satisfies the $2^{p-1}$-approximate triangle inequality [^6] Here, $p$ can be a real number., and assume $\mathbb{D}$ to be uniform distribution. We show that in balanced markets with adversarial servers and stochastic requests, an $O(1)$-competitive algorithm exists for $d\geq 1$ and $p\in[1,\max\{d,2\})$.
@@ B.1 Proof of Theorem 8
Appendix C An Alternative Algorithm for Tree Metrics
In this section, we consider balanced markets and apply our approach to tree metrics with general distributions. In particular, a tree metric is defined by shortest path distances in a tree $T=(V,E)$. We will prove that the $\mathsf{SOAR}$ algorithm admits an $O(1)$ competitive ratio with both servers and requests randomly drawn from $\mathbb{D}$, which implies the existence of an $O(1)$-competitive algorithm with servers adversarially chosen and requests randomly drawn from $\mathbb{D}$ by Theorem 1. This gives an alternative $O(1)$-competitive algorithm in balanced markets with tree metrics and general distributions, recovering the result by [GGPW19].
@@ Appendix C An Alternative Algorithm for Tree Metrics
Suppose that $m=n$ and consider tree metrics. For an arbitrary distribution $\mathbb{D}$, an $O(1)$-competitive algorithm exists when servers are adversarially chosen and requests are randomly drawn from $\mathbb{D}$.
@@
According to Lemma 15, we can assume requests to be uniformly distributed over server locations by losing a constant factor in the competitive ratio. In other words, let $a_{v}$ be the number of servers located at $v\in V$, the distribution $\mathbb{D}$ satisfies $\mathbb{P}_{r\sim\mathbb{D}}[r=v]=a_{v}/n$ for every $v\in V$. By Theorem 1, it suffices to give an $O(1)$-competitive algorithm when both servers and requests are randomly drawn from $\mathbb{D}$.
Let ${\mathcal{A}}$ be the algorithm as per Lemma 5. Recall that $\mathrm{OPT}(t)$ is estimated by Lemma 16. When both servers and requests are randomly drawn from $\mathbb{D}$, the expected cost of ${\mathcal{A}}$ equals

### v2

@@ Appendix C Stochastic Setting: Cost Functions Beyond Triangle Inequality
The proof of Theorem 10 can be obtained by replacing every usage of the triangle inequality by the approximate triangle inequality in the proofs of Theorem 1 and Theorem 2 and hence is omitted.
Next, we consider the case where $X=[0,1]^{d}$ and $\delta(x,y)=\left\|{x-y}\right\|_{2}^{p}$ for $p\geq 1$, [^6] Here, $p$ can be a real number. which we will prove that satisfies the $2^{p-1}$-approximate triangle inequality, and assume $\mathbb{D}$ to be uniform distribution. We show that in balanced markets with adversarial servers and stochastic requests, an $O(1)$-competitive algorithm exists for $d\geq 1$ and $p\in[1,\max\{d,2\})$.
@@ C.1 Proof of Theorem 11
Appendix D Stochastic Setting: An Alternative Algorithm for Tree Metrics
In this section, we consider balanced markets and apply our approach to tree metrics with general distributions. In particular, a tree metric is defined by shortest path distances in a tree $T=(V,E)$. We will prove that the $\mathsf{SOAR}$ algorithm admits an $O(1)$ competitive ratio with all servers and requests independently drawn from $\mathbb{D}$, which implies the existence of an $O(1)$-competitive algorithm with servers adversarially chosen and requests independently drawn from $\mathbb{D}$ by Theorem 2. This gives an alternative $O(1)$-competitive algorithm in balanced markets with tree metrics and general distributions, recovering the result by [GGPW19].
@@ Appendix D Stochastic Setting: An Alternative Algorithm for Tree Metrics
Suppose that $m=n$ and consider tree metrics. For an arbitrary distribution $\mathbb{D}$, an $O(1)$-competitive algorithm exists when servers are adversarially chosen and requests are independently drawn from $\mathbb{D}$.
@@
According to Lemma 18, we can assume requests to be uniformly distributed over server locations by losing a constant factor in the competitive ratio. In other words, let $a_{v}$ be the number of servers located at $v\in V$, the distribution $\mathbb{D}$ satisfies $\mathbb{P}_{r\sim\mathbb{D}}[r=v]=a_{v}/n$ for every $v\in V$. By Theorem 2, it suffices to give an $O(1)$-competitive algorithm when all servers and requests are independently drawn from $\mathbb{D}$.
Let ${\mathcal{A}}$ be the algorithm as per Lemma 5. Recall that $\mathrm{OPT}(t)$ is estimated by Lemma 19. When all servers and requests are independently drawn from $\mathbb{D}$, the expected cost of ${\mathcal{A}}$ equals